An underwater diver sees the sun 50° above horizontal. How high is the sun above the horizon to a fisherman in a boat above the diver?
Answer #1

from snell's law of refraction \(n_{\text {water }} \sin \theta_{\text {water }}=n_{\text {air }} \sin \theta_{\text {air }}\)

\(\theta_{\text {water }}=90-50=40\)

\(n_{\text {water }}=1.33\)

\(n_{\text {air }}=1\)

\(1.33 \times \sin 40=1 \times \sin \theta_{\text {air }}\)

\(\theta_{\text {air }}=58.7\)

the fisherman in boat can see at an angle \(=90-58.7=31.3^{\circ}\)